5V- 46 23 5 7

A wire is connected to each side of Bulb 2, shown above. Calculate the new equivalent resistance if each bulb has a 30 ohm resistance.

a) 95 ohms

b) 90 ohms

c) 110 ohms

or d) 85 ohms
Answer #1

when a wire is connected across the bulb it means it is short circuited and R2=0.

so bulb2 and bulb 3 are in series so their equivalent resistance is

R'=R1+R2=30+0=30 ohm

Now R' and R4 will be in parallel

so R''=R'*R4/(R'+R4)=30*30/(30+30)=900/60=15 ohm

Now bulb 5 and bulb6 are in parallel so

R'''=R5*R6/(R5+R6)=30*30/(30+30)=900/60=15 ohm

now R''' and R'' are in series so

R''''=R'''+R''=15+15=30 ohm

Now at last bulb1 and bulb7 and R'''' are in series so equivalent resistance

R_{eq}=30+30+30=90ohm